# SENSE models: an open source solution for multilingual and multimodal semantic-based tasks

### v1

@@ Abstract
Index Terms:
multilingual speech encoder, semantic representation, multimodal information retrieval, speech translation
@@ I Introduction
Speech foundation models based on self-supervised learning (SSL) have brought significant advances in speech processing. These models, such as wav2vec 2.0 [1], HuBERT [2], and WavLM [3], generate learned speech representations that can be applied to a wide range of downstream speech processing tasks. By training on large amounts of unlabelled speech data, SSL models have demonstrated the ability to capture crucial speech features, such as phonemes and other acoustic units [4]. This capability has led to significant progress in multiple downstream tasks, including speech recognition [1], speech translation [5], speech separation, speaker verification, speaker diarization [3], and emotion detection [6].
@@
Different approaches have been proposed to pretrain model by aligning speech and text, like mSLAM [7], a Massively multilingual joint pre-training for speech and text. In 2022, MIT and LIUM introduced a novel approach to learn semantically-aligned, multimodal, utterance-level, cross-lingual speech representations (SAMU-XLSR: Semantically-Aligned Multimodal Utterance-level cross-lingual speech representation)[8] from SSL pretrained speech encoders. The embedding vector space of SAMU-XLSR is both multimodal and cross-lingual: it is shared between speech and text modalities and is common to multiple languages. Moreover, it is semantically aligned, as spoken utterances are clustered together with their corresponding speech and text translations within the vector space. To achieve this alignment, [8] proposes a multimodal learning framework that fine-tunes the pre-trained multilingual XLSR-128 speech encoder [9] through a teacher-student approach led by the pre-trained language-agnostic BERT sentence encoder, LaBSE [10]. More recently, META AI has also presented the SONAR approach [11] as a part of the Seamless models for speech translation [12]. The SAMU-XLSR and SONAR approaches are very similar since they both consist of making the models able to extract and represent semantics at the utterance level by learning from a textual model (respectively LaBSE [10] and a specialized NLLB decoder [13]). They allow reaching very good results on tasks that require the extraction of semantics from speech (speech translation, spoken language understanding, speech retrieval…). While this approach is very promising, there is no open-source solution that could help researchers aiming to work in this direction. There is no public implementation of SAMU-XLSR, and even if the SONAR models are distributed under an open-source licence and the source code to use them, the code to retrain similar models is not available.
@@
This paper introduces SENSE (Shared Embedding for N-lingual Speech and tExt), an open-source solution directly inspired by the SAMU-XLSR framework. We describe how the original SAMU-XLSR method has been updated by selecting a stronger teacher text model and a better initial speech encoder. The source code for training and using SENSE models has been integrated into the SpeechBrain toolkit [14], and the first SENSE model we trained has been publicly released [^1] https://huggingface.co/LIA-AvignonUniversity. For reproducibility reasons, we release the first SENSE model that have been trained using publically-available data. We report experimental results on multilingual and multimodal semantic tasks, where our initial SENSE model achieves highly competitive performance. Finally, this study offers new insights into how semantics are captured in such multilingual, semantically aligned speech encoders.
@@ II SAMU-XSLR framework
SAMU-XLSR has been introduced [8] in order to fine-tune the multilingual XLS-R SSL speech encoder [9]. Frame-level representations are merged by an attentive pooling layer, then passed through a linear projection and tanh activation function, producing one utterance-level vector per audio file while keeping XLS-R weights adaptable (Fig. 1).
This speech vector is trained to imitate the LaBSE text embedding of the utterance transcript [10]. The cosine similarity between the LaBSE text embedding and the utterance-level SAMU-XSLR is maximized through the loss function. Because LaBSE supplies a language-agnostic space covering 109 languages, SAMU-XLSR learns cross-lingual speech–text alignment by simply pulling each speech vector toward its multilingual text anchor, with no explicit cross-lingual audio supervision.
@@ III-A Teacher model selection
The Massive Text Embedding Benchmark (MTEB) [15] aims to provide a comprehensive evaluation of text embedding models across a wide range of tasks. It includes 58 datasets spanning 112 languages and covers 8 different embedding tasks. We used the MTEB leaderboard [^2] https://huggingface.co/spaces/mteb/leaderboard to identify the model that best meets our specific requirements. Among the top-performing models, BGE-M3 [16] stood out due to its strong performance in various tasks such as reranking, bitext mining, and semantic textual similarity. In addition to its performance, BGE-M3 is relatively lightweight with 560 million parameters, fully open-source, and multilingual, supporting over 100 languages. It can also handle documents of varying lengths, with support for sequences of up to 8,192 tokens.
@@ III-B SSL speech encoder selection
The ML-SUPERB Challenge (Multilingual Speech Universal PERformance Benchmark) [17] provides a large-scale benchmark for evaluating multilingual speech processing systems, covering many languages. Based on the 2025 baseline results  [^3] https://multilingual.superbbenchmark.org/challenge-interspeech2025/challenge_overview trained using the ML-SUPERB 1-hour dataset, we observe that w2v-BERT 2.0 outperforms other models with fewer than 1 billion parameters. A benchmark on Tunisian Arabic [18] has shown that the best open-source speech encoder is w2v-BERT 2.0. Internal benchmarks were conducted on different languages and across various acoustic conditions (clean and noisy data) and confirmed this finding.
@@ III-C SENSE training details
We developed our own implementation in SpeechBrain to build models of the SAMU-XLSR type. As mentioned above, we selected a new teacher model, BGE-M3 [16], and a new speech encoder model w2v-BERT 2.0 [19] to initialize the SENSE model. We then trained it on the publicly available Common Voice 19 dataset [20], using only the languages supported by the BGE-M3 model, a total of 83 languages, corresponding to 8,250 hours of speech. The Common Voice dataset is organized into two sets: validated and unvalidated. The validated partition consists of audio samples that have been reviewed and confirmed by human annotators to have accurate transcriptions and good audio quality. Within the validated set, the data is further divided into train, development, and test subsets. To train the SENSE model, we used only the validated training partition. Details regarding the languages and the duration of each language subset are provided in the appendix -A.
Following [8], we perform either up- or down-sampling based on each language’s representation in the dataset. We initialized our speech encoder model with the w2v-BERT 2.0 model. The w2v-BERT 2.0 model comprises 24 Conformer layers [21] and combines contrastive and masked prediction learning and is pre-trained on 4.5 million hours of unlabelled audio data that covers 143 languages. In order to build our fixed-size utterance representation, we added an attention pooling [22] operation on the encoder outputs. This model is pre-trained for 100K iterations by employing 32 H100 GPUs. We select the batch size to optimize GPU memory usage, resulting in 20 min of audio per batch. We use different learning rates and optimizers for the pre-trained encoder and the attention pooling. We fine-tune the speech encoder using a learning rate of $10^{-5}$ while the attention pooling is trained with a learning rate of 1. We used the cosine distance between the speech and the text embedding as the training loss. The text encoder is kept frozen throughout the training.
@@ V-A Multilingual and Multimodal Translation Retrieval task
To perform translation retrieval, we construct two databases (DB): a query database and a search database. The query database contains speech utterances in language X, while the search database consists of either speech or text in language Y, depending on the scenario. We consider many scenarios: speech $\to$ speech retrieval and speech $\to$ text retrieval. The objective is to retrieve the correct translation in a different language. Depending on the scenario, each audio segment or text sentence in the query database is transformed into an embedding. If the modality is speech, a speech encoder is used to extract the embedding at the utterance level. If the modality is text, a text encoder is applied to extract the sentence embedding. Similarly, the entries in the search database are processed in the same way, using the appropriate encoder based on their modality. Once both sets of data (query database and search database) are turned into embeddings that can be compared and normalized via mean subtraction, translation retrieval is performed by computing the similarity between the query embeddings and all entries in the search database. Following [8], we evaluate the translation retrieval task using Recall@1.
@@ X Speech $\to$ EN Speech retrieval
We use the public VoxPopuli dataset [23], which provides a large-scale multilingual corpus suitable for speech-to-speech translation and other tasks. We follow [8] and use the same languages. This setup involves speech queries in one language and a search DB composed of their corresponding translations in English. The query DB and the search DB are equal in size.
@@ X Speech $\to$ Y Speech retrieval
We select a subset of random language pairs from the VoxPopuli [23] dataset used previously, where $X\neq Y$, $X\neq\text{EN}$, and $Y\neq\text{EN}$.
@@
We use the publicly available MUST-C [24] dataset. MUST-C is a large-scale multilingual corpus designed specifically for speech-to-text translation tasks. The speech query DB is in English and the search DB is in another language. We follow [8] to create the query and search databases.
@@ X Speech $\to$ EN Text
We use the public MTEDx dataset [25]. In order to evaluate the generalization ability of the models to unseen languages, we additionally used the FLEURS dataset [26].
@@ X Speech $\to$ Y Text
We use the FLEURS dataset [26] and MTEDx dataset [25] by selecting a subset of language pairs.
@@ V-B Spoken Language Understanding tasks
We evaluate our model on MSNER [27], a multilingual NER benchmark based on subsets of the VoxPopuli dataset in four languages (Dutch, French, German, and Spanish). The dataset is annotated using the 18 entity classes defined in OntoNotes [28], providing a structured and linguistically rich evaluation setting. We use the official partitions of the dataset.
@@
For the slot filling task, we use SpeechMassive [29] SpeechMassive is a dataset designed to evaluate SLU capabilities in both monolingual and cross-lingual settings. The annotation format in SpeechMassive follows the SLURP schema [30]. In total, the dataset defines 55 distinct slot types. We used also the SLURP dataset. Since for each task, we include some evaluation for low-resource languages not covered by the learned model, we evaluate SLU also using TARIC-SLU [31]. The TARIC-SLU dataset is a publicly available resource designed to advance SLU in Tunisian Arabic. It comprises over 2,000 dialogues. These dialogues are annotated with 62 semantic slot types.
The study in [32] demonstrated the effectiveness of the semantically aware frame-level speech representations of SAMU-XLSR for a specific semantic extraction task. In line with this work, we adopt a frame-level approach in our evaluation. We formulate the SLU (for both NER and Slot filling) task as a character level prediction where slots or named entities are delimited by tag specific special characters, as in [33, 34]. As input, the neural network receives an audio file and the output is a transcription enriched with semantic labels. In addition to the speech encoder model, we incorporate an extra layer with 1024 neurons and LeakyReLU as the activation function, followed by a fully-connected layer and a final softmax layer where each dimension corresponding to a character. The weights of these two additional layers were randomly initialized, while the weights of the speech encoder part for SSL models of the neural architecture were initialized using the pretrained weights. We evaluate the NER and slot filling tasks using Named Entity Error Rate (NEER) [34], COncept Error Rate (COER) and Concept Value Error Rate (CVER) [35]. COER and NEER are computed similarly to Word Error Rate by taking into account only the semantic (or entities) labels in the reference and hypothesis annotations. The CVER computation is identical, but the occurrences of concept/value pairs are taken into account instead of the concept alone.
@@ V-C Speech summarization
For Speech summarization, we use DECODA dataset [36]. DECODA is a French human-human spoken conversation dataset composed of spoken dialogues. Each conversation is accompanied by a manual transcription and a short summary that captures the key points discussed in the conversation. Building upon the work of [37], we employ the same version of the DECODA dataset. In this task, we use a large-scale pretrained sequence-to-sequence model called BARThez [38], specifically designed for French. We trained the BARThez model to generate summaries from speech embeddings. The first training stage uses the French MLSUM dataset, as it is one of the largest summarization datasets available in French. This step is essential to help the model adapt to sentence embeddings as input. The second training stage specializes the model on task-specific dataset. We fine-tune the model on DECODA using speech utterance embeddings, to generate summaries directly from speech. This task is evaluated using ROUGE-L [39] and BertScore [40].
@@ V-D Speech-To-Text translation
We evaluated speech-to-text translation task using the CoVoST 2 dataset [41]. CoVoST 2 is derived from Common Voice [20] with additionnal translations.
We use the standard encoder-decoder architecture for our translation model. We initialize the encoder with a speech encoder and decoder using mBart [42]. A feed-forward network projection layer is used to connect the encoder and decoder, bridging the two modules. BLEU is used to evaluate the translation task.
@@ VI Experimental results
This section presents the experimental results on all tasks described above. We compare SENSE with two state-of-the-art models, SAMU-XLSR and SONAR, for experiments based on utterance-level embeddings. For frame-level experiments, the comparison is made with w2v-BERT 2.0 and SAMU-XLSR. The SAMU-XLSR used in our work is the version trained on 53 languages, following the same training procedure as the 25-language version presented in [8]. SENSE and SAMU-XLSR use a single multilingual speech encoder shared across all languages. SONAR, on the other hand, uses 37 language-specific encoders. In all our experiments, we use the corresponding speech encoder when the language is available; otherwise, we use the English speech encoder.

### v2

@@ I Introduction
Speech foundation models based on self-supervised learning (SSL) have brought significant advances in speech processing. These models, such as wav2vec 2.0 [baevski2020wav2vec], HuBERT [hsu2021hubert], and WavLM [chen2022wavlm], generate learned speech representations that can be applied to a wide range of downstream speech processing tasks. By training on large amounts of unlabelled speech data, SSL models have demonstrated the ability to capture crucial speech features, such as phonemes and other acoustic units [pasad2023comparative]. This capability has led to significant progress in multiple downstream tasks, including speech recognition [baevski2020wav2vec], speech translation [nguyen2020investigating], speech separation, speaker verification, speaker diarization [chen2022wavlm], and emotion detection [macary2021use].
@@
Different approaches have been proposed to pretrain model by aligning speech and text, like mSLAM [bapna2022mslam], a Massively multilingual joint pre-training for speech and text. In 2022, MIT and LIUM introduced a novel approach to learn semantically-aligned, multimodal, utterance-level, cross-lingual speech representations (SAMU-XLSR: Semantically-Aligned Multimodal Utterance-level cross-lingual speech representation)[khurana2022samu] from SSL pretrained speech encoders. The embedding vector space of SAMU-XLSR is both multimodal and cross-lingual: it is shared between speech and text modalities and is common to multiple languages. Moreover, it is semantically aligned, as spoken utterances are clustered together with their corresponding speech and text translations within the vector space. To achieve this alignment, [khurana2022samu] proposes a multimodal learning framework that fine-tunes the pre-trained multilingual XLSR-128 speech encoder [babu2021xls] through a teacher-student approach led by the pre-trained language-agnostic BERT sentence encoder, LaBSE [feng2022language]. More recently, META AI has also presented the SONAR approach [Duquenne:2023:sonar_arxiv] as a part of the Seamless models for speech translation [seamless2025joint]. The SAMU-XLSR and SONAR approaches are very similar since they both consist of making the models able to extract and represent semantics at the utterance level by learning from a textual model (respectively LaBSE [feng2022language] and a specialized NLLB decoder [nllbteam2022languageleftbehindscaling]). They allow reaching very good results on tasks that require the extraction of semantics from speech (speech translation, spoken language understanding, speech retrieval…). While this approach is very promising, there is no open-source solution that could help researchers aiming to work in this direction. There is no public implementation of SAMU-XLSR, and even if the SONAR models are distributed under an open-source licence and the source code to use them, the code to retrain similar models is not available.
@@
This paper introduces SENSE (Shared Embedding for N-lingual Speech and tExt), an open-source solution directly inspired by the SAMU-XLSR framework. We describe how the original SAMU-XLSR method has been updated by selecting a stronger teacher text model and a better initial speech encoder. The source code for training and using SENSE models has been integrated into the SpeechBrain toolkit [ravanelli2024open], and the first SENSE model we trained has been publicly released [^1] https://huggingface.co/LIA-AvignonUniversity. For reproducibility reasons, we release the first SENSE model that have been trained using publically-available data. We report experimental results on multilingual and multimodal semantic tasks, where our initial SENSE model achieves highly competitive performance. Finally, this study offers new insights into how semantics are captured in such multilingual, semantically aligned speech encoders.
@@ II SAMU-XSLR framework
SAMU-XLSR has been introduced [khurana2022samu] in order to fine-tune the multilingual XLS-R SSL speech encoder [babu2021xls]. Frame-level representations are merged by an attentive pooling layer, then passed through a linear projection and tanh activation function, producing one utterance-level vector per audio file while keeping XLS-R weights adaptable (Fig. 1).
This speech vector is trained to imitate the LaBSE text embedding of the utterance transcript [feng2022language]. The cosine similarity between the LaBSE text embedding and the utterance-level SAMU-XSLR is maximized through the loss function. Because LaBSE supplies a language-agnostic space covering 109 languages, SAMU-XLSR learns cross-lingual speech–text alignment by simply pulling each speech vector toward its multilingual text anchor, with no explicit cross-lingual audio supervision.
@@ III-A Teacher model selection
The Massive Text Embedding Benchmark (MTEB) [muennighoff2023mteb] aims to provide a comprehensive evaluation of text embedding models across a wide range of tasks. It includes 58 datasets spanning 112 languages and covers 8 different embedding tasks. We used the MTEB leaderboard [^2] https://huggingface.co/spaces/mteb/leaderboard to identify the model that best meets our specific requirements. Among the top-performing models, BGE-M3 [chen2024m3] stood out due to its strong performance in various tasks such as reranking, bitext mining, and semantic textual similarity. In addition to its performance, BGE-M3 is relatively lightweight with 560 million parameters, fully open-source, and multilingual, supporting over 100 languages. It can also handle documents of varying lengths, with support for sequences of up to 8,192 tokens.
@@ III-B SSL speech encoder selection
The ML-SUPERB Challenge (Multilingual Speech Universal PERformance Benchmark) [yang2024mlsuperb] provides a large-scale benchmark for evaluating multilingual speech processing systems, covering many languages. Based on the 2025 baseline results  [^3] https://multilingual.superbbenchmark.org/challenge-interspeech2025/challenge_overview trained using the ML-SUPERB 1-hour dataset, we observe that w2v-BERT 2.0 outperforms other models with fewer than 1 billion parameters. A benchmark on Tunisian Arabic [mdhaffar2024performance] has shown that the best open-source speech encoder is w2v-BERT 2.0. Internal benchmarks were conducted on different languages and across various acoustic conditions (clean and noisy data) and confirmed this finding.
@@ III-C SENSE training details
We developed our own implementation in SpeechBrain to build models of the SAMU-XLSR type. As mentioned above, we selected a new teacher model, BGE-M3 [chen2024m3], and a new speech encoder model w2v-BERT 2.0 [barrault2023seamless] to initialize the SENSE model. We then trained it on the publicly available Common Voice 19 dataset [ardila2020common], using only the languages supported by the BGE-M3 model, a total of 83 languages, corresponding to 8,250 hours of speech. The Common Voice dataset is organized into two sets: validated and unvalidated. The validated partition consists of audio samples that have been reviewed and confirmed by human annotators to have accurate transcriptions and good audio quality. Within the validated set, the data is further divided into train, development, and test subsets. To train the SENSE model, we used only the validated training partition. Details regarding the languages and the duration of each language subset are provided in the appendix -A. Following [khurana2022samu], we perform either up- or down-sampling based on each language’s representation in the dataset. We initialized our speech encoder model with the w2v-BERT 2.0 model. The w2v-BERT 2.0 model comprises 24 Conformer layers [gulati2020conformer] and combines contrastive and masked prediction learning and is pre-trained on 4.5 million hours of unlabelled audio data that covers 143 languages. In order to build our fixed-size utterance representation, we added an attention pooling [santos2016attentive] operation on the encoder outputs. This model is pre-trained for 100K iterations by employing 32 H100 GPUs. We select the batch size to optimize GPU memory usage, resulting in 20 min of audio per batch. We use different learning rates and optimizers for the pre-trained encoder and the attention pooling. We fine-tune the speech encoder using a learning rate of $10^{-5}$ while the attention pooling is trained with a learning rate of 1. We used the cosine distance between the speech and the text embedding as the training loss. The text encoder is kept frozen throughout the training.
@@ V-A Multilingual and Multimodal Translation Retrieval task
To perform translation retrieval, we construct two databases (DB): a query database and a search database. The query database contains speech utterances in language X, while the search database consists of either speech or text in language Y, depending on the scenario. We consider many scenarios: speech $\to$ speech retrieval and speech $\to$ text retrieval. The objective is to retrieve the correct translation in a different language. Depending on the scenario, each audio segment or text sentence in the query database is transformed into an embedding. If the modality is speech, a speech encoder is used to extract the embedding at the utterance level. If the modality is text, a text encoder is applied to extract the sentence embedding. Similarly, the entries in the search database are processed in the same way, using the appropriate encoder based on their modality. Once both sets of data (query database and search database) are turned into embeddings that can be compared and normalized via mean subtraction, translation retrieval is performed by computing the similarity between the query embeddings and all entries in the search database. Following [khurana2022samu], we evaluate the translation retrieval task using Recall@1.
@@ X Speech $\to$ EN Speech retrieval
We use the public VoxPopuli dataset [wang2021voxpopuli], which provides a large-scale multilingual corpus suitable for speech-to-speech translation and other tasks. We follow [khurana2022samu] and use the same languages. This setup involves speech queries in one language and a search DB composed of their corresponding translations in English. The query DB and the search DB are equal in size.
@@ X Speech $\to$ Y Speech retrieval
We select a subset of random language pairs from the VoxPopuli [wang2021voxpopuli] dataset used previously, where $X\neq Y$, $X\neq\text{EN}$, and $Y\neq\text{EN}$.
@@
We use the publicly available MUST-C [di2019must] dataset. MUST-C is a large-scale multilingual corpus designed specifically for speech-to-text translation tasks. The speech query DB is in English and the search DB is in another language. We follow [khurana2022samu] to create the query and search databases.
@@ X Speech $\to$ EN Text
We use the public MTEDx dataset [salesky2021multilingual]. In order to evaluate the generalization ability of the models to unseen languages, we additionally used the FLEURS dataset [conneau2023fleurs].
@@ X Speech $\to$ Y Text
We use the FLEURS dataset [conneau2023fleurs] and MTEDx dataset [salesky2021multilingual] by selecting a subset of language pairs.
@@ V-B Spoken Language Understanding tasks
We evaluate our model on MSNER [meeus2024msner], a multilingual NER benchmark based on subsets of the VoxPopuli dataset in four languages (Dutch, French, German, and Spanish). The dataset is annotated using the 18 entity classes defined in OntoNotes [weischedel11272ontonotes], providing a structured and linguistically rich evaluation setting. We use the official partitions of the dataset.
@@
For the slot filling task, we use SpeechMassive [lee2024speech] SpeechMassive is a dataset designed to evaluate SLU capabilities in both monolingual and cross-lingual settings. The annotation format in SpeechMassive follows the SLURP schema [bastianelli2020slurp]. In total, the dataset defines 55 distinct slot types. We used also the SLURP dataset. Since for each task, we include some evaluation for low-resource languages not covered by the learned model, we evaluate SLU also using TARIC-SLU [mdhaffar2024taric]. The TARIC-SLU dataset is a publicly available resource designed to advance SLU in Tunisian Arabic. It comprises over 2K dialogues. These dialogues are annotated with 62 semantic slot types. We used also MEDIA dataset [laperriere2022spoken].
The study in [laperriere2023use] demonstrated the effectiveness of the semantically aware frame-level speech representations of SAMU-XLSR for a specific semantic extraction task. In line with this work, we adopt a frame-level approach in our evaluation. We formulate the SLU (for both NER and Slot filling) task as a character level prediction where slots or named entities are delimited by tag specific special characters, as in [ghannay2018end, mdhaffar2022end]. As input, the neural network receives an audio file and the output is a transcription enriched with semantic labels. In addition to the speech encoder model, we incorporate an extra layer with 1024 neurons and LeakyReLU as the activation function, followed by a fully-connected layer and a final softmax layer where each dimension corresponding to a character. The weights of these two additional layers were randomly initialized, while the weights of the speech encoder part for SSL models of the neural architecture were initialized using the pretrained weights. We evaluate the NER and slot filling tasks using Named Entity Error Rate (NEER) [mdhaffar2022end], COncept Error Rate (COER) and Concept Value Error Rate (CVER) [laperriere2022spoken]. COER and NEER are computed similarly to Word Error Rate by taking into account only the semantic (or entities) labels in the reference and hypothesis annotations. The CVER computation is identical, but the occurrences of concept/value pairs are taken into account instead of the concept alone.
@@ V-C Speech summarization
For Speech summarization, we use DECODA dataset [bechet2012decoda]. DECODA is a French human-human spoken conversation dataset composed of spoken dialogues. Each conversation is accompanied by a manual transcription and a short summary that captures the key points discussed in the conversation. Building upon the work of [akani2024unified], we employ the same version of the DECODA dataset. In this task, we use a large-scale pretrained sequence-to-sequence model called BARThez [eddine2021barthez], specifically designed for French. We trained the BARThez model to generate summaries from speech embeddings. The first training stage uses the French MLSUM dataset, as it is one of the largest summarization datasets available in French. This step is essential to help the model adapt to sentence embeddings as input. The second training stage specializes the model on task-specific dataset. We fine-tune the model on DECODA using speech utterance embeddings, to generate summaries directly from speech. This task is evaluated using ROUGE-L [lin2004rouge] and BertScore [zhang2019bertscore].
@@ V-D Speech-To-Text translation
We evaluated speech-to-text translation task using the CoVoST 2 dataset [wang2021covost]. CoVoST 2 is derived from Common Voice [ardila2020common] with additionnal translations.
We use the standard encoder-decoder architecture for our translation model. We initialize the encoder with a speech encoder and decoder using mBart [liu2020mbart]. A feed-forward network projection layer is used to connect the encoder and decoder, bridging the two modules. BLEU is used to evaluate the translation task.
@@ VI Experimental results
This section presents the experimental results on all tasks described above. We compare SENSE with two state-of-the-art models, SAMU-XLSR and SONAR, for experiments based on utterance-level embeddings. For frame-level experiments, the comparison is made with w2v-BERT 2.0 and SAMU-XLSR. The SAMU-XLSR used in our work is the version trained on 53 languages, following the same training procedure as the 25-language version presented in [khurana2022samu]. SENSE and SAMU-XLSR use a single multilingual speech encoder shared across all languages. SONAR, on the other hand, uses 37 language-specific encoders. In all our experiments, we use the corresponding speech encoder when the language is available; otherwise, we use the English speech encoder.